OpenStudy (anonymous):

1. Factor x2 + 64. Check your work.

OpenStudy (anonymous):

@ranga

OpenStudy (anonymous):

@Adriennne

OpenStudy (anonymous):

please help?!

OpenStudy (ranga):

a^2 + b^2 = (a + ib) * (a - ib) where i is square root of -1. x^2 + 64 = x^2 + 8^2 Apply the formula in line 1 to line 2.

OpenStudy (anonymous):

I know it has something to do with complex numbers.... ok, so (x+8i)(x-8i)??

OpenStudy (anonymous):

is that right?

OpenStudy (ranga):

Yes.

OpenStudy (anonymous):

Ok!!! Thank you soo much !!!!

OpenStudy (ranga):

You are welcome.

OpenStudy (anonymous):

:)

OpenStudy (ranga):

They say to check your work. So multiply (x+8i)(x-8i) using the FOIL method and verify the answer is indeed x^2 + 64.